OpenStudy (anonymous):

The surface area of a cube is 864 in2. What is the area of one face of this cube?

rishavraj (rishavraj):

surface area of cube = 6a^2 so \[6 a^2 = 864\] once u get the side length i.e "a" u can calculate the area of one face

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

what 6*6*6=216 not 864

OpenStudy (anonymous):

i dont get it

rishavraj (rishavraj):

see the surface area of a cube with side length "a" = \[6 a^2\] the given surface area is 864 so \[6 a^2 = 864\]

OpenStudy (anonymous):

so 6 is the ans

OpenStudy (wil476003):

No ur over thinking it you just have to divide 864 by 6 so the answer is 144

OpenStudy (anonymous):

o ok

OpenStudy (mathstudent55):

A cube is made up of 6 faces. You are given the total surface ares of the 6 faces. You want the surface area of one single face. This is a one-step problem involving only one operation.

rishavraj (rishavraj):

no no .....once u solve tht ....i mean get the side length